You believe the price of Rose, Inc., stock is going to fall and you’ve decided to sell 900 shares short. If the current share price is $53, construct the equity account balance sheet for this trade. Assume the initial margin is 100 percent. (Input all amounts as positive values.)  

Assets Liabilities and account equity
Proceeds from sale Short position
Initial margin deposit Account equity
Total $0 Total $0
Answer #1

Proceeds from sale = Number of shares sell short* the current share price = 900*53= $47,700

Initial Margin= Proceeds from sale* 1.00= $47,700*1= $47,700

Assets
Liabilities and account equity
Proceeds from sale
$47,700
Short position
$47,700
Initial margin deposit
$47,700
Account equity
$47,700
Total $95,400 Total $95,400